The area of a rectangle is 35 yd2, and the length of the rectangle is
8 yd less than three times the width. Find the dimensions of the rectangle.

Answer by Boreal(15235)

x=width
3x-8=length
3x^2-8x=35
3x^2-8x-35=0
(3x+7)(x-5)=0
x=5, and x=-7/3, which we cannot use.
width is 5 yds.
length is 7 yds.